rootbeer003:

A small auditorium has 10 rows of seats. There are 12 seats in the first row and 16 seats in the second row. The number of seats in a row continues to increase by 4 with each additional row. What is the total number of seats in the auditorium? 352 156 300 252

rootbeer003:

@Hero

Hero:

@Ultrilliam can probably answer this question

Ultrilliam:

I'm busy ATM dealing with ASO and their poor decisions

Hero:

@rootbeer003, how many questions do you have in total?

rootbeer003:

3

Hero:

You have three more questions left?

rootbeer003:

including this one yes

Hero:

Okay so this is what you do. We know there are ten rows of seats so we "count" 1st row --- 12 seats 2nd row --- 16 seats 3rd row --- 20 seats ... nth row --- a_n seats So we have to figure out a relationship between n and a_n. We do this by starting with the first row and counting in the following manner: 1st row --- 12 + 0(4) seats 2nd row --- 12 + 1(4) seats 3rd row --- 12 + 2(4) seats ... nth row --- 12 + (n -1)(4) seats

Hero:

We use this pattern to figure out the number of seats in the 10th row.

Hero:

10th row ---> 12 + (10 - 1)(4) seats So @rootbeer003 how many seats are there in the 10th row?

rootbeer003:

48

Hero:

Very good. Now we have to figure out the total number of seats in the auditorium. Do you have the sum formula with you?

rootbeer003:

no

Hero:

\(\sum_{\text{n = 1}}^{10} 12 + (n-1)(4)\)

Hero:

Ugh, latex is terrible with sums

rootbeer003:

idk how to use that formula..

Hero:

@Ultrilliam you should fix this. LaTeX doesn't work well with Sum formulas at all

rootbeer003:

is it 300?

Hero:

Are you guessing or did you actually use a formula to come up with that?

Hero:

The other way to figure it out is to use the following formula: \(\text{Sum} = n\left(\dfrac{a_1 + a_n}{2}\right)\)

rootbeer003:

What do i put in as a and n

Hero:

In this case, \(n = 10, a_1 = 12, a_n = 48\), \(\text{Sum} = n\left(\dfrac{a_1 + a_n}{2}\right)\)

Hero:

In other words n represents total number of rows \(a_1\) represents the first term in the sequence \(a_n\) represents the last term in the sequence

Hero:

The answer is 300, but you did not explain how you came up with it.

rootbeer003:

1 attachment
Hero:

Post as a separate question please.